# hep-th/yymmnnn

### v1

@@ Abstract
In this paper, we study black hole solutions in double field theory. In the first part, we introduce a solution-generating method and classify black hole solutions into three categories in standard double field theory. To solve the problem of double time, we utilize space/time split double field theory in the second part to derive black hole solutions. By introducing a cosmological constant and imposing the strong constraint, our findings indicate that when considering the entire doubled spacetime, the curvature signifies an AdS vacuum. However, in the subspace of the entire spacetime, black holes and curvature singularities emerge.
@@ 1 Introduction
$S=\int d^{4}x\sqrt{-g}e^{-2\phi}\left[R+4\left(\nabla\phi\right)^{2}\right].$ (1.1)
@@
2nd kind of solutions: If $g_{ij}\left(x\right)$ and $\left(g^{-1}\right)_{ij}\left(x\right)$ are self-dual solutions of the effective action (1.1), the constraint-satisfying solutions of DFT can be given as:
@@
3rd kind of solutions: If $g_{ij}\left(x\right)$ and $\left(g^{-1}\right)_{ij}\left(x\right)$ are self-dual solutions of the effective action (1.1), we can construct constraint-violating solutions of DFT such that $g\left(\tilde{x},x\right)=g\left(\tilde{x}\right)\times\left(g^{-1}\right)\left(x\right)$. These solutions depend on double coordinates simultaneously and take the forms:
@@
The problem with these solutions is the presence of double timelike directions $dt$ and $d\tilde{t}$, which lack explicit physical interpretations. To solve this issue, we consider the $12D-$ 1 dimensional space/time split DFT, which is independent of the dual time coordinate. To obtain non-trivial solutions, we consider non-critical strings, which introduce the cosmological constant. We then calculate the equations of motion (EOM) for the corresponding action and use the generated methods mentioned above to derive the black hole solutions. The new insights of these solutions arise from the curvature definition in doubled spacetime. In DFT, the Ricci curvature is generalized to include the doubled coordinates and dilatonic dynamics, known as the generalized Ricci scalar $\mathcal{R}$. It satisfies the EOM for the dilaton and always vanishes ($\mathcal{R}=0$). Since we consider the cosmological constant, the generalized Ricci scalar for the whole doubled spacetime is:
@@
This implies that, regardless of the obtained solution, the entire doubled spacetime is an AdS vacuum. On the other hand, since the geometry of the subspace $x$ of the doubled spacetime can be completely determined by the low-energy effective action, we can construct the black hole solution in this subspace. In other words, considering the entire spacetime results in an AdS vacuum, whereas black holes and curvature singularities emerge in the subspace.
@@ 2.1 Equations of motion for the generalized metric and dilaton
In summary, the dilatonic and gravitational EOM of DFT are presented as follows:
@@ 2.3 Classification of black hole solutions
If $g_{ij}\left(x\right)$ and $\phi\left(x\right)$ are solutions of the action (2.48), and $\left(g^{-1}\right)_{ij}\left(x\right)$ is also a solution for the same action, these configurations are termed self-dual solutions. They satisfy the strong constraint and encompass four distinct types, expressed as:
@@
To illustrate this self-dual property, consider the following ansatz:
@@ A. Generalized scalar curvature
indicating that our solutions represent AdS vacuum solutions in double coordinates. Importantly, DFT possesses AdS vacuum solutions due to the generalized Ricci scalar includes dilatonic effects.
@@ B. Reduction to low-energy effective action
This entire metric describes the AdS vacuum in double coordinates. The first bracket corresponds to the black hole solution known from ordinary string theory, while the second bracket represents the additional dimensions introduced by DFT. Since the first bracket is fully determined by the low-energy effective action, the usual interpretations in this framework can be directly applied. This suggests that this conventional, undoubled spacetime $x$ is self-contained and exhibits its own dynamics. If we consider the additional part $\left(\tilde{r},\tilde{\varphi}\right)$, there is no action to describe its geometry independently. Therefore, we must consider the entire doubled spacetime. DFT governs its dynamics, and it manifests as an AdS vacuum. In other words, if we obscure the doubled coordinates and focus only on the conventional region, the black hole and its singularity emerge as solutions of the low-energy effective action, as depicted in Figure. (1).
@@
Figure 1: The dark blue region represents the conventional undoubled spacetime described by the low-energy effective action, which contains the black holes. When the additional region from the double coordinates is included, only DFT can describe the entire doubled spacetime, consistently characterizing it as an AdS vacuum.
@@ 1. When $a=1$, $b=0$, $c=1$, we obtain:
If we only consider the first square bracket, the corresponding Kretschmann scalar is $R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}=\frac{\lambda^{4}M^{2}}{r^{2}}$, indicating this black hole solution possesses a spacetime singularity: $r=0$ where the event horizon is located at $r=M$. Considering the whole metric, there are two time-like directions $dr^{2}$ and $d\tilde{r}^{2}$ acrossing the event horizon. Additionally, accounting for the coordinates $\left(\tilde{r},\tilde{\varphi}\right)$, there is a coordinate singularity as $r\rightarrow\infty$. However, it is impossible to measure the curvature in the region $\left(\tilde{r},\tilde{\varphi}\right)$ independently. When considering the doubled coordinates, the metric corresponds to the AdS vacuum.
@@ 2. When $a=-1$, $b=0$, $c=1$, we have:
Considering the first square bracket, it is the solution of low-energy effective action and the corresponding Kretschmann scalar is $R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}=\frac{\lambda^{4}M^{2}}{\left(M-r\right)^{2}}$. This solution has a single curvature singularity at $r=M$, with a coordinate singularity as $r\rightarrow\infty$. And for complete metric, the metric still corresponds to the AdS vacuum.
@@ 3. When $a=0$, $b=\pm 1$, $c=1$, we have:
Here, $b=1$ for $M\geq 0$ and $b=-1$ for $M\leq 0$. When $b=-1$, there is a conical singularity at $r=M$ due to $R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}=\frac{\lambda^{4}M^{2}}{\left(M-r\right)^{2}}$. When $b=1$, the curvature singularity is at $r=0$ because $R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}=\frac{\lambda^{4}M^{2}}{r^{2}}$. Furthermore, when the doubled coordinates are considered, the metric corresponds to the AdS vacuum.
@@ 4. When $a=b=c=0$, the solution is given as follows:
The first square bracket of this metric describe flat spacetime, as $R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}=0$. This implies that the half-geometry of the AdS vacuum in the doubled geometry can also manifest as flat spacetime when the doubled coordinates are ignored.
@@ 4 Conclusion and discussion
In this paper, we studied black hole solutions in DFT across two main sections. Initially, we directly calculated black hole solutions within standard DFT, where metrics involve two time-like directions. These solutions were categorized into three classes: two imposing to the strong constraint and one violating it by depending on dual coordinates. Due to the ambiguity in interpreting solutions with two time-like directions, further investigation is warranted for these cases. In the second part, we explored the space/time split DFT, where the theory only depends on conventional time-like directions, manifesting in a $1+2\left(D-1\right)$ dimensional spacetime $\left(t,x,\tilde{x}\right)$. In this framework, we investigated non-critical string theory incorporating a cosmological constant, also deriving solutions that violate the constraint. These solutions, similarly, are left for future investigation. Furthermore, we derived double black string solutions satisfying the strong constraint. In the context of the entire doubled spacetime, the generalized Ricci scalar consistently manifests as a negative constant, indicating AdS vacuum solutions. Upon hidden the dual coordinates $x$, the standard geometry described by $\tilde{x}$ reveals black hole solutions and curvature singularities.
@@
Acknowledgements We are deeply indebt to Houwen Wu for many discussions and suggestions. This work was supported by NSFC Grant No.12105031, No.12347101 and PSFC Grant No. cstc2021jcyj-bshX0227.

### v3

@@ Abstract
In this paper, we study black hole solutions in double field theory. In the first part, we introduce a solution-generating method and classify black hole solutions into three categories in standard double field theory. To solve the problem of double time, we utilize space/time split double field theory in the second part to derive black hole solutions. By introducing a cosmological constant and imposing the strong constraint, our findings indicate that when considering the entire doubled spacetime, the curvature signifies a generalized AdS vacuum. However, in the subspace of the entire spacetime, black holes and curvature singularities emerge.
@@ 1 Introduction
$S=\int d^{D}x\sqrt{-g}e^{-2\phi}\left[R+4\left(\nabla\phi\right)^{2}\right].$ (1.1)
@@
2nd kind of solutions: If $g_{ij}\left(x\right)$ and $\left(g^{-1}\right)_{ij}\left(x\right)$ are solutions of the effective action (1.1), the constraint-satisfying solutions of DFT can be given as [^1] This kind of solutions are also discussed in Refs. [38, 39].:
@@
3rd kind of solutions: If $g_{ij}\left(x\right)$ and $\left(g^{-1}\right)_{ij}\left(x\right)$ are solutions of the effective action (1.1), we can construct constraint-violating solutions of DFT such that $g\left(\tilde{x},x\right)=g\left(\tilde{x}\right)\times\left(g^{-1}\right)\left(x\right)$. These solutions depend on double coordinates simultaneously and take the forms:
@@
The problem with these solutions is the presence of double timelike directions $dt$ and $d\tilde{t}$, which lack explicit physical interpretations. To solve this issue, we consider the $12D-$ 1 dimensional space/time split DFT, which is independent of the dual time coordinate. To obtain non-trivial solutions, we consider non-critical strings, which introduce the cosmological constant. This additional term is valid as the DFT action should reduce to the low-energy effective action after imposing the strong constraint [6, 8]. We then calculate the equations of motion (EOM) for the corresponding action and use the generated methods mentioned above to derive the black hole solutions. The new insights of these solutions arise from the curvature definition in doubled spacetime. In DFT, the Ricci curvature is generalized to include the doubled coordinates and dilatonic dynamics, known as the generalized Ricci scalar $\mathcal{R}$. It satisfies the EOM for the dilaton and always vanishes ($\mathcal{R}=0$). Since we consider the cosmological constant, the generalized Ricci scalar for the whole doubled spacetime is:
@@
This implies that, regardless of the obtained solution, the entire doubled spacetime is a generalized AdS vacuum. On the other hand, since the geometry of the subspace $x$ of the doubled spacetime can be completely determined by the low-energy effective action, we can construct the black hole solution in this subspace. In other words, considering the entire spacetime results in the generalized AdS vacuum, whereas black holes and curvature singularities emerge in the subspace. Finally, we note that a complete understanding of DFT black holes is still far from being achieved. The primary challenge is the lack of evidence confirming whether the generalized metric can be interpreted as the physical spacetime of our universe. Therefore, our focus is on obtaining solutions to the DFT action and providing potential interpretations.
@@ 2.1 Equations of motion for the generalized metric and dilaton
This is a T-duality covariant generalized tensor, which serves as a natural extension of the Ricci tensors in Riemannian geometry. In summary, the dilatonic and gravitational EOM of DFT are presented as follows:
@@ 2.3 Classification of black hole solutions
If $g_{ij}\left(x\right)$ and $\phi\left(x\right)$ are solutions of the action (2.48), and $\left(g^{-1}\right)_{ij}\left(x\right)$ is also a solution for the same action. They satisfy the strong constraint and encompass four distinct types, expressed as:
@@
To illustrate this property, consider the following ansatz:
@@ A. Generalized scalar curvature
indicating that our solutions represent the generalized AdS vacuum solutions in double coordinates. This additional term is valid as the DFT action should reduce to the non-critical low-energy effective action after imposing the strong constraint [6, 8]. Importantly, DFT possesses the generalized AdS vacuum solutions due to the generalized Ricci scalar includes dilatonic effects. However, we wish to note that this solution differs from the traditional AdS vacuum, as it is derived from the generalized Ricci scalar, which is defined on the doubled spacetime. Further studies are needed to fully understand its physical implications.
@@ B. Reduction to low-energy effective action
This entire metric describes the generalized AdS vacuum in double coordinates. The first bracket corresponds to the black hole solution known from ordinary string theory, while the second bracket represents the additional dimensions introduced by DFT. Since the first bracket is fully determined by the low-energy effective action, the usual interpretations in this framework can be directly applied. This suggests that this conventional, undoubled spacetime $x$ is self-contained and exhibits its own dynamics. If we consider the additional part $\left(\tilde{r},\tilde{\varphi}\right)$, there is no action to describe its geometry independently. Therefore, we must consider the entire doubled spacetime. DFT governs its dynamics, and it manifests as the generalized AdS vacuum. In other words, if we obscure the doubled coordinates and focus only on the conventional region, the black hole and its singularity emerge as solutions of the low-energy effective action, as depicted in Figure. (1).
@@
Figure 1: The dark blue region represents the conventional undoubled spacetime described by the low-energy effective action, which contains the black holes. When the additional region from the double coordinates is included, only DFT can describe the entire doubled spacetime, consistently characterizing it as the generalized AdS vacuum.
@@ 1. When $a=1$, $b=0$, $c=1$, we obtain:
If we only consider the first square bracket, the corresponding Kretschmann scalar is $R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}=\frac{\lambda^{4}M^{2}}{r^{2}}$, indicating this black hole solution possesses a spacetime singularity: $r=0$ where the event horizon is located at $r=M$. Considering the whole metric, there are two time-like directions $dr^{2}$ and $d\tilde{r}^{2}$ acrossing the event horizon. Additionally, accounting for the coordinates $\left(\tilde{r},\tilde{\varphi}\right)$, there is a coordinate singularity as $r\rightarrow\infty$. However, it is impossible to measure the curvature in the region $\left(\tilde{r},\tilde{\varphi}\right)$ independently. When considering the doubled coordinates, the metric corresponds to the generalized AdS vacuum.
@@ 2. When $a=-1$, $b=0$, $c=1$, we have:
Considering the first square bracket, it is the solution of low-energy effective action and the corresponding Kretschmann scalar is $R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}=\frac{\lambda^{4}M^{2}}{\left(M-r\right)^{2}}$. This solution has a single curvature singularity at $r=M$, with a coordinate singularity as $r\rightarrow\infty$. And for complete metric, the metric still corresponds to the generalized AdS vacuum.
@@ 3. When $a=0$, $b=\pm 1$, $c=1$, we have:
Here, $b=1$ for $M\geq 0$ and $b=-1$ for $M\leq 0$. When $b=-1$, there is a conical singularity at $r=M$ due to $R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}=\frac{\lambda^{4}M^{2}}{\left(M-r\right)^{2}}$. When $b=1$, the curvature singularity is at $r=0$ because $R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}=\frac{\lambda^{4}M^{2}}{r^{2}}$. Furthermore, when the doubled coordinates are considered, the metric corresponds to the generalized AdS vacuum.
@@ 4. When $a=b=c=0$, the solution is given as follows:
The first square bracket of this metric describe flat spacetime, as $R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}=0$. This implies that the half-geometry of the generalized AdS vacuum in the doubled geometry can also manifest as flat spacetime when the doubled coordinates are ignored.
Comment on the Cosmological Solutions
Finally, we wish to note that our calculation methods for DFT black holes are also applicable to DFT cosmology. We start with the low-energy effective action, assuming a vanishing Kalb-Ramond field:
$S=\int d^{D}x\sqrt{-g}e^{-2\phi}\left[R+4\left(\nabla\phi\right)^{2}\right].$ (3.117)
The cosmological solutions of DFT can be constructed through the solutions of this action (3.117):
1st kind of solutions: If $g_{ij}\left(t\right)$ and $\phi\left(t\right)$ are solutions of the effective action (3.117), the solutions of DFT that satisfy the strong constraint can be given as:
$ds^{2}=g^{ij}\left(t\right)d\tilde{x}_{i}d\tilde{x}_{j}+g_{ij}\left(t\right)dx^{i}dx^{j},\qquad or\qquad ds^{2}=\left(g^{-1}\right)^{ij}\left(\tilde{t}\right)d\tilde{x}_{i}d\tilde{x}_{j}+\left(g^{-1}\right)_{ij}\left(\tilde{t}\right)dx^{i}dx^{j}.$ (3.118)
Note this case works for any cosmological solution of (3.117).
2nd kind of solutions: If $g_{ij}\left(t\right)$ and $\left(g^{-1}\right)_{ij}\left(t\right)$ are $O\left(d,d\right)$ dual solutions of the effective action (3.117), the constraint-satisfying solutions of DFT can be given as:
$ds^{2}=g^{ij}\left(t\right)d\tilde{x}_{i}d\tilde{x}_{j}+g_{ij}\left(t\right)dx^{i}dx^{j},\qquad or\qquad ds^{2}=\left(g^{-1}\right)^{ij}\left(t\right)d\tilde{x}_{i}d\tilde{x}_{j}+\left(g^{-1}\right)_{ij}\left(t\right)dx^{i}dx^{j},$ (3.119)
or
$ds^{2}=g^{ij}\left(\tilde{t}\right)d\tilde{x}_{i}d\tilde{x}_{j}+g_{ij}\left(\tilde{t}\right)dx^{i}dx^{j},\qquad or\qquad ds^{2}=\left(g^{-1}\right)^{ij}\left(\tilde{t}\right)d\tilde{x}_{i}d\tilde{x}_{j}+\left(g^{-1}\right)_{ij}\left(\tilde{t}\right)dx^{i}dx^{j}.$ (3.120)
For example, the simplest solutions are given by:
$ds^{2}=-d\tilde{t}^{2}+\left|t\right|^{\pm 1/\sqrt{D-1}}\delta_{ij}d\tilde{x}_{i}d\tilde{x}_{j}-dt^{2}+\left|t\right|^{\mp 1/\sqrt{D-1}}\delta_{ij}dx^{i}dx^{j},\qquad d\left(t\right)=-\frac{1}{2}\ln\left|t\right|,$ (3.121)
or
$ds^{2}=-d\tilde{t}^{2}+\left|\tilde{t}\right|^{\pm 1/\sqrt{D-1}}\delta_{ij}d\tilde{x}_{i}d\tilde{x}_{j}-dt^{2}+\left|\tilde{t}\right|^{\mp 1/\sqrt{D-1}}\delta_{ij}dx^{i}dx^{j},\qquad d\left(\tilde{t}\right)=-\frac{1}{2}\ln\left|\tilde{t}\right|,$ (3.122)
where $d\left(t\right)$ and $d\left(\tilde{t}\right)$ are $O\left(d,d\right)$ dilaton.
3rd kind of solutions: If $g_{ij}\left(t\right)$ and $\left(g^{-1}\right)_{ij}\left(t\right)$ are solutions of the effective action (3.117), we can construct constraint-violating solutions of DFT such that $g\left(\tilde{t},t\right)=g\left(\tilde{t}\right)\times\left(g^{-1}\right)\left(t\right)$. These solutions depend on double time coordinates and take the forms:
$ds^{2}=g^{ij}\left(t,\tilde{t}\right)d\tilde{x}_{i}d\tilde{x}_{j}+g_{ij}\left(t,\tilde{t}\right)dx^{i}dx^{j},\qquad or\qquad ds^{2}=\left(g^{-1}\right)^{ij}\left(t,\tilde{t}\right)d\tilde{x}_{i}d\tilde{x}_{j}+\left(g^{-1}\right)_{ij}\left(t,\tilde{t}\right)dx^{i}dx^{j}.$ (3.123)
For example, the simplest solutions are given by:
$ds^{2}=-d\tilde{t}^{2}+\left|\frac{t}{\tilde{t}}\right|^{\pm 1/\sqrt{D-1}}\delta_{ij}d\tilde{x}_{i}d\tilde{x}_{j}-dt^{2}+\left|\frac{t}{\tilde{t}}\right|^{\mp 1/\sqrt{D-1}}\delta_{ij}dx^{i}dx^{j},\qquad d\left(t,\tilde{t}\right)=-\frac{1}{2}\ln\left|tt\right|,$ (3.124)
or
$ds^{2}=-d\tilde{t}^{2}+\left|t\tilde{t}\right|^{\pm 1/\sqrt{D-1}}\delta_{ij}d\tilde{x}_{i}d\tilde{x}_{j}-dt^{2}+\left|t\tilde{t}\right|^{\mp 1/\sqrt{D-1}}\delta_{ij}dx^{i}dx^{j},\qquad d\left(t,\tilde{t}\right)=-\frac{1}{2}\ln\left|t\tilde{t}\right|.$ (3.125)
@@ 4 Conclusion and discussion
In this paper, we studied black hole solutions in DFT across two main sections. Initially, we directly calculated black hole solutions within standard DFT, where metrics involve two time-like directions. These solutions were categorized into three classes: two imposing to the strong constraint and one violating it by depending on dual coordinates. Due to the ambiguity in interpreting solutions with two time-like directions, further investigation is warranted for these cases. In the second part, we explored the space/time split DFT, where the theory only depends on conventional time-like directions, manifesting in a $1+2\left(D-1\right)$ dimensional spacetime $\left(t,x,\tilde{x}\right)$. In this framework, we investigated non-critical string theory incorporating a cosmological constant, also deriving solutions that violate the constraint. These solutions, similarly, are left for future investigation. Furthermore, we derived double black string solutions satisfying the strong constraint. In the context of the entire doubled spacetime, the generalized Ricci scalar consistently manifests as a negative constant, indicating the generalized AdS vacuum solutions. Upon hidden the dual coordinates $x$, the standard geometry described by $\tilde{x}$ reveals black hole solutions and curvature singularities.
@@
Acknowledgements This work was supported by NSFC Grant No.12105031, No.12347101 and PSFC Grant No. cstc2021jcyj-bshX0227.